OpenStudy (anonymous):

Evaluate the integral. 0~3(6+6y-y^2)dy a.-18 b.45 c.54 d.36 e.-12

OpenStudy (mimi_x3):

\[ \int 6+6y-y^{2} dy\]?

OpenStudy (anonymous):

yupp with 0 at the bottom and 3 at the top

OpenStudy (mimi_x3):

Where are you having problems integrating?

OpenStudy (anonymous):

\[\large = \int\limits_{0}^{3}6.dy + \int\limits_{0}^{3}6y.dy - \int\limits_{0}^{3}y^2.dy\]

OpenStudy (anonymous):

the whole thing for this one i just need the answer because we already got them in class i just lost the paper and then we are supposed to do something with the answers

OpenStudy (anonymous):

36

OpenStudy (anonymous):

I have more if you can help

OpenStudy (anonymous):

yep

OpenStudy (anonymous):

\[\int\limits_{-2}^{5}\left| 4x-x ^{2} \right|dx\]

OpenStudy (anonymous):

\[\int\limits_{0}^{1}3x ^{2}\cos(x ^{3})dx\]

OpenStudy (anonymous):

do you know how to start the standard integrals?

OpenStudy (anonymous):

\[if \int\limits_{0}^{6} f(x)dx=10 and \int\limits_{0}^{4} f(x)dx= 7 find \int\limits_{4}^{6}f(x)dx\]

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ans is 3 for first one

OpenStudy (anonymous):

that isnt a choice

OpenStudy (anonymous):

109, 123, 102, 209, 23